# CHD: Coupled Hierarchical Diffusion for Long-Horizon Tasks

### v2

@@ 1 Introduction
Diffusion models have achieved strong results in image generation [1], video synthesis [2], and protein modeling [3]. Recently, they’ve been applied to robot control, where diffusion-based planners generate smooth, coherent, and multi-modal trajectories [4, 5]. However, as planning horizons grow, trajectory variance increases, uncertainty compounds, and achieving high reward becomes more difficult [6]. Hierarchical diffusion planners offer a promising approach towards addressing this problem by decomposing planning into high-level (HL) subgoal inference and low-level (LL) goal-conditioned trajectory generation [7, 8]—this decomposition reduces horizon length and distribution complexity.
@@ 2.1 Problem Formulation
We focus on finite-horizon, maximum-reward trajectory optimization in a discrete-time system [4]. Let $s_{k}\!\in\!\mathcal{S}$ and $a_{k}\!\in\!\mathcal{A}$ denote the state and action at time step $k=0,\dots,H$, with dynamics $s_{k+1}=f(s_{k},a_{k})$. During planning, we aim to generate a trajectory $\bm{\tau}=(s_{0},a_{0},\dots,s_{H},a_{H})$ that maximizes the cumulative reward $\mathcal{J}(\bm{\tau})=\sum_{k=0}^{H}r(s_{k},a_{k})$.
@@ 2.2 Diffusion Models as Planners
Diffusion probabilistic models [1] are generative models that synthesize data by iteratively denoising noisy samples from $T$ to 00. They can also be used to generate trajectories $\bm{\tau}$ [4], where the forward process $q(\bm{\tau}_{t}|\bm{\tau}_{t-1})$ progressively corrupts the data (“clean” trajectories), while the reverse process $p_{\theta}(\bm{\tau}_{t-1}|\bm{\tau}_{t})$ denoises the data. The induced marginal distribution over trajectories is $p_{\theta}(\bm{\tau}_{0})=\int p(\bm{\tau}_{T})\prod_{t=1}^{T}p_{\theta}(\bm{\tau}_{t-1}|\bm{\tau}_{t})d\bm{\tau}_{1:T},$
@@ 1. P1:
    Bi-directional Coupling. High-level (HL) subgoals guide low-level (LL) trajectory generation, and LL feedback refines HL subgoals.
@@ 3.1 Joint Diffusion Model (JDM)
JDM is a principled and expressive formulation that naturally provides bidirectional coupling (P1) between high-level and low-level components. However, because it diffuses over the entire joint space, it does not reduce the planning horizon or data complexity (violating P3), nor does it allow independent, parallel sampling of HL and LL (violating P2), which limits its practical scalability.
@@ 3.2 Coupled Hierarchical Diffusion Planner
Removing the direct dependence of the HL kernel on the LL state yields a more modular structure that we augment below with classifier guidance, asynchronous sampling, and segment-wise planning to satisfy P1–P3.
@@ 4.1 Maze Navigation
We begin with the Maze Navigation benchmark [10], where trajectories lie in continuous 2D space and HL subgoals correspond to key intermediate joint states between LL segments. The agent receives a reward of $+1$ upon reaching the target and 00 otherwise; thus, the optimal policy reaches the goal in the fewest steps. This task is challenging due to sparse rewards and sub-optimal demonstration data. We compare CHD against baseline diffusion planners, including Diffuser [4], Decision Diffuser (DD) [5] with classifier-free guidance, and baseline hierarchical diffusion methods (BHD), HMDI (BHD with graph-search improved subgoals) [9] and SHD (BHD with evenly divided subgoals) [7].
@@
Environment Diffuser DD BHD HMDI SHD CHD (ours)
Maze2D U-Maze $113.9\pm 3.1$ $116.2\pm 2.7$ $118.5\pm 5.4$ $120.1\pm 2.5$ $128.4\pm 3.6$ $\textbf{142.8}\pm 2.9$
@@ 4.3 Analysis and Discussion
CHD improves sampling efficiency through parallel generation. We benchmarked training and inference times on 8$\times$×NVIDIA RTX A5000 GPUs (Table 3). CHD and BHD both train significantly faster than flat Diffuser models; the hierarhical approach reduces distribution complexity, which allowed us to reduce the number of diffusion steps from 256 to 32 while maintaining performance. However, at inference time, BHD requires sequential generation of HL and LL trajectories, which increases sampling time. In contrast, CHD supports asynchronous parallel generation on GPUs by coupling HL and LL updates within the diffusion process. This leads to a 30–40% reduction in sampling time compared to BHD, while maintaining superior planning performance.
@@
Methods Training ($h$)1 Sampling ($s$)2
@@ D.5 Algorithm
1:  $p\left(\bm{\tau}_{T}^{g}\right)\sim\mathcal{N}\left(\bm{\tau}_{T}^{g};\mathbf{0},\mathbf{I}^{x}\right)$, $p\left(\bm{\tau}_{T}^{x}\right)\sim\mathcal{N}\left(\bm{\tau}_{T}^{x};\mathbf{0},\mathbf{I}^{g}\right)$
@@
7:  Calculate $p_{\theta^{x}}(\bm{\tau}_{0}^{x}|\bm{\tau}_{0}^{g},\bm{\tau}_{1}^{x})$ and apply LL classifier guidance $p_{\phi}(\mathcal{O}_{1:N}=1|\bm{\tau}^{g}_{0},\bm{\tau}^{x}_{0})$
@@ D.7 Advantages of CHD over BHD
[…]
@@
[…]
@@ E.1 Maze Navigation Experiments
Figure 8: Maze navigation environment in D4RL. The agent start position (greed dot) is randomly chosen in the maze, and the goal position (red dot) is either fixed (Maze2D) or also randomized (Multi2D). The agent will get a “$+1$” reward when reaching the goal.
@@
Maze navigation is a long-horizon trajectory planning environment in the D4RL dataset (Maze2D)  [10]. The environment has three maze configurations with different sizes: U-maze, Medium and Large. In the simulator, the agent is initially placed in the maze and the target is to plan a feasible trajectory to reach the goal position. For each maze configuration, D4RL can either set a fixed goal position called “Maze2D” or a random goal position in the maze called “Multi2D”, increasing the tasks’ variance and requiring higher generalization ability (Fig. 8). In addition, the agent gets a “$+1$” reward when it reaches the vicinity of the goal position. So in a maze navigation environment, the planned trajectory should not only successfully reach the goal, but also reduce the time spent traveling to earn higher cumulated rewards.
@@
We finally evaluate the goal-reaching performance using the default method in D4RL. When the agent reaches the goal, it gets a “$+1$” reward and we calculate the overall reward in one episode. Then we applied maximin normalization and ×$\times 100$ to compare the results. Higher values mean that the trajectory used fewer steps to reach the goal. We run the experiments with 150 different seeds and report and mean and standard deviation of the results.
@@
Episode length Segment2 interval Planner horizon Diffusion steps
Single1 HL2 LL2 Single HL LL
@@
[…]
@@ E.2 Robot Task Planning Experiments
The robot task planning experiments were conducted using the Kitchen World simulator [13], which is designed for learning-based algorithms in task planning [11, 13] and supports integration with downstream solvers like PDDLStream [62]. In Kitchen World, the PR2 dual-arm robot moves in the kitchen scenario, manipulating rigid and articulated objects to finish certain tasks. In our experiments, we created a customized task-planning benchmark with extended horizons to evaluate performance. During the training phase, we utilized offline demonstrations to train the planners. In the sampling phase, the planner generated sub-goal states and actions, which were subsequently executed using the PDDLStream solver as the policy.
[…]
@@
We constructed the dataset around the task of “cooking meals” within the kitchen environment, comprising 10 sub-tasks. Fig. 12 illustrates an example scene and demonstration. The state and action spaces, as well as initial states, are predefined in the scene. The program then generates HL subgoal states and LL actions required to accomplish the target task. The order of the 10 sub-tasks is interchangeable, provided they collectively achieve the target, which enhances the model’s generalization capability in real-world scenarios.
To diversify demonstrations, we randomized the initial and terminal states of objects, varied the sub-task sequences, and introduced redundant behaviors. This approach ensured suboptimal demonstrations with extended trajectories. Based on the average trajectory lengths in the demonstrations, tasks were categorized into Easy (50 steps), Medium (70 steps), and Hard (90 steps). Additionally, a subset of tasks with fixed terminal states was defined as “Single,” while the full dataset with randomized terminal states was labeled “Multi.” The dataset consisted of $10,000$ training samples and $1,000$ validation samples. Finally, we evaluated the planned tasks in the simulator by measuring the number of completed tasks and the steps required to execute them.
@@
Average demo. length Episode length Segment interval Planner horizon Diffusion steps
Single1 HL2 LL2 Single HL LL
@@
We elaborate on implementing Bits Diffusion [16] in the task planning experiments. In Fig. 14, we first defined the whole state and action spaces, which were tokenized using the PreTrainedTokenizerFast in Huggingface (code). Then we converted the tokens as 8 Bits analog values. Since the state and action have two tokens, the diffusion model generates values with dimension “horizon ×$\times 16$”. In the sampling stage, the planners sampled values in continuous space. And we converted them with thresholding to binary values. Finally, the bits were digitalized and de-tokenized back to the texts for the following steps.

### v3

@@ 1 Introduction
Diffusion models have achieved strong results in image generation [1], video synthesis [2], and protein modeling [3]. Recently, they’ve been applied to robot control, where diffusion-based planners generate smooth, coherent, and multi-modal trajectories [4, 5, 6, 7]. However, as planning horizons grow, trajectory variance increases, uncertainty compounds, and achieving high reward becomes more difficult [8]. Hierarchical diffusion planners offer a promising approach towards addressing this problem by decomposing planning into high-level (HL) subgoal inference and low-level (LL) goal-conditioned trajectory generation [9, 10]—this decomposition reduces horizon length and distribution complexity.
@@ 2.1 Problem Formulation
We focus on finite‑horizon, maximum‑reward trajectory optimization in a discrete‑time system [4]. Let $s_{k}\!\in\!\mathcal{S}$ and $a_{k}\!\in\!\mathcal{A}$ denote the state and action at time step $k=0,\dots,H$, with dynamics $s_{k+1}=f(s_{k},a_{k})$. During planning, we aim to generate a trajectory $\bm{\tau}=(s_{0},a_{0},\dots,s_{H},a_{H})$ that maximizes the cumulative reward $\mathcal{J}(\bm{\tau})=\sum_{k=0}^{H}r(s_{k},a_{k})$.
@@ 2.2 Diffusion Models as Planners
Diffusion probabilistic models [1] are generative models that synthesize data by iteratively denoising noisy samples from $T$ to 0. They can also be used to generate trajectories $\bm{\tau}$ [4], where the forward process $q(\bm{\tau}_{t}|\bm{\tau}_{t-1})$ progressively corrupts the data (“clean” trajectories), while the reverse process $p_{\theta}(\bm{\tau}_{t-1}|\bm{\tau}_{t})$ denoises the data. The induced marginal distribution over trajectories is $p_{\theta}(\bm{\tau}_{0})=\int p(\bm{\tau}_{T})\prod_{t=1}^{T}p_{\theta}(\bm{\tau}_{t-1}|\bm{\tau}_{t})d\bm{\tau}_{1:T},$
@@ 1. P1:
    Bi-directional Coupling. High‑level (HL) subgoals guide low‑level (LL) trajectory generation, and LL feedback refines HL subgoals.
@@ 3.1 Joint Diffusion Model (JDM)
JDM is a principled and expressive formulation that naturally provides bidirectional coupling (P1) between high‑level and low‑level components. However, because it diffuses over the entire joint space, it does not reduce the planning horizon or data complexity (violating P3), nor does it allow independent, parallel sampling of HL and LL (violating P2), which limits its practical scalability. Therefore, evaluating JDM is equivalent to the flat diffusion planner.
@@ 3.2 Coupled Hierarchical Diffusion Planner
Removing the direct dependence of the HL kernel on the LL state yields a more modular structure that we augment below with classifier guidance, asynchronous sampling, and segment‑wise planning to satisfy P1–P3.
@@ 4.1 Maze Navigation
We begin with the Maze Navigation benchmark [12], where trajectories lie in continuous 2D space and HL subgoals correspond to key intermediate joint states between LL segments. The agent receives a reward of +1+1 upon reaching the target and 0 otherwise; thus, the optimal policy reaches the goal in the fewest steps. This task is challenging due to sparse rewards and sub-optimal demonstration data. We compare CHD against baseline diffusion planners, including Diffuser [4], Decision Diffuser (DD) [5] with classifier-free guidance, and baseline hierarchical diffusion methods (BHD), HMDI (BHD with graph-search improved subgoals) [11] and SHD (BHD with evenly divided subgoals) [9].
@@
Environment Diffuser $D$ BHD HMDI SHD CHD (ours)
Maze2D U-Maze $13.9\pm 3.1 11$$6.2\pm 2.7 11$$18.5\pm 5.4$ $120.1\pm 2.5$ $128.4\pm 3.6$ $\textbf{142.8}\pm 2.9$
@@ 4.3 Analysis and Discussion
[…]
@@
Methods Training (hh)1 Sampling (ss)2
@@ 5 Conclusion
Acknowledgment. This research/project is supported by A*STAR under its National Robotics Programme (NRP) (Award M23NBK0053) and the Ministry of Education, Singapore, under the Academic Research Fund Tier 1 (T1 251RES2515). The authors would also like to acknowledge support from Google.
@@ D.5 Algorithm
[…]
@@
7: Calculate $p_{\theta^{x}}(\bm{\tau}_{0}^{x}|\bm{\tau}_{0}^{g},\bm{\tau}_{1}^{x})$ and apply $L$ classifier guidance $p_{\phi}(\mathcal{O}_{1:N}=1|\bm{\tau}^{g}_{0},\bm{\tau}^{x}_{0})$
@@ D.7 Advantages of CHD over BHD
$\text{BHD:}\quad p(\bm{\tau}^{x},\bm{\tau}^{g}\mid O=1)=p(\bm{\tau}^{g})p_{\theta^{x}}(\bm{\tau}^{x}\mid\bm{\tau}^{g})p(O=1\mid\bm{\tau}^{g})$ (37)
$\text{CHD:}\quad p(\bm{\tau}^{x},\bm{\tau}^{g}\mid O=1)=p(\bm{\tau}^{g})p(\bm{\tau}^{x}\mid\bm{\tau}^{g})p(O=1\mid{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bm{\tau}^{x}},\bm{\tau}^{g})$ (38)
@@
$\int p(O=1\mid\tilde{\bm{\tau}}^{x},\bm{\tau}^{g})p(\tilde{\bm{\tau}}^{x}\mid\bm{\tau}^{g})d\tilde{\bm{\tau}}^{x}\leq\int p(O=1\mid\bm{\tau}^{x},\bm{\tau}^{g})p(\tilde{\bm{\tau}}^{x}\mid\bm{\tau}^{g})d\tilde{\bm{\tau}}^{x}$ (41)
Also, we know the integration of a probability is smaller than 1 $\int p(\tilde{\bm{\tau}}^{x}\mid\bm{\tau}^{g})d\tilde{\bm{\tau}}^{x}\leq 1$. Bring this to the inequality and we finally prove that CHD achieves higher optimality than BHD. $p(O=1\mid\bm{\tau}^{g})\leq p(O=1\mid\bm{\tau}^{x},\bm{\tau}^{g})$. Equality holds when the optimality $O$ is independent from HL subgoals $\bm{\tau}^{x}$.
@@ E.1 Maze Navigation Experiments
Figure 8: Maze navigation environment in D4RL. The agent start position (greed dot) is randomly chosen in the maze, and the goal position (red dot) is either fixed (Maze2D) or also randomized (Multi2D). The agent will get a “+1+1” reward when reaching the goal.
@@
Maze navigation is a long-horizon trajectory planning environment in the D4RL dataset (Maze2D)  [12]. The environment has three maze configurations with different sizes: U-maze, Medium and Large. In the simulator, the agent is initially placed in the maze and the target is to plan a feasible trajectory to reach the goal position. For each maze configuration, D4RL can either set a fixed goal position called “Maze2D” or a random goal position in the maze called “Multi2D”, increasing the tasks’ variance and requiring higher generalization ability (Fig. 8). In addition, the agent gets a “+1+1” reward when it reaches the vicinity of the goal position. So in a maze navigation environment, the planned trajectory should not only successfully reach the goal, but also reduce the time spent traveling to earn higher cumulated rewards.
@@
We finally evaluate the goal-reaching performance using the default method in D4RL. When the agent reaches the goal, it gets a “+1+1” reward and we calculate the overall reward in one episode. Then we applied maximin normalization and $\times 100$ to compare the results. Higher values mean that the trajectory used fewer steps to reach the goal. We run the experiments with 150 different seeds and report and mean and standard deviation of the results.
@@
Episode
length
Segment2
interval
Planner horizon Diffusion steps
Single1 HL2 LL2 Single HL $L$
@@
We present the hyperparameters used in our implementation in Table 5. For Diffuser, we directly adopted the hyperparameters from the official code. Then, we adapted the hyperparameters in BHD and CHD to achieve a hierarchical structure. A notable modification was the reduction of diffusion steps, as the data distribution for shorter horizons is less complex and therefore easier to model. Additionally, the segment interval is defined as $\text{LL horizon}-1$, since the endpoints correspond to two adjacent LL segments.
@@ E.2 Robot Task Planning Experiments
The robot task planning experiments were conducted using the Kitchen World simulator [15], which is designed for learning-based algorithms in task planning [13, 15] and supports integration with downstream solvers like PDDLStream [64]. In Kitchen World, the PR2 dual-arm robot moves in the kitchen scenario, manipulating rigid and articulated objects to finish certain tasks. In our experiments, we created a customized task-planning benchmark with extended horizons to evaluate performance. During the training phase, we utilized offline demonstrations to train the planners. In the sampling phase, the planner generated sub-goal states and actions, which were subsequently executed using the PDDLStream solver as the policy. In short, the simplified PDDL maps the semantic action and targets to the robot and objects in the kitchen world simulation and plans trajectories to satisfy each primitive motion.
Task planning was formalized in a simplified PDDL space [8]. The state space is represented as $s=[(\texttt{Object},\texttt{Position})_{n}]$, where $n$ is the number of objects, and the action space is defined as $a=[(\texttt{Motion},\texttt{Target})]$. The Object set includes all movable rigid and articulated objects in the environment, such as $\texttt{Bowl},\texttt{Pot},\texttt{Cabinet},\dots]$. The Position specifies the location or state of the object, e.g., $\texttt{On-table},\texttt{In-cabinet},\texttt{Open},\texttt{In-Bowl},\dots]$. The Motion component of actions describes the robot’s behavior at each step, encompassing $\texttt{Pick},\texttt{Place},\texttt{Grasp},\texttt{Push},\texttt{Move},\dots]$. The Target specifies the intended target of the motion, such as $\texttt{Chicken},\texttt{Cabinet},\texttt{Right-Table},\dots]$.
@@
We constructed the dataset around the task of “cooking meals” within the kitchen environment, comprising 10 sub-tasks. Fig. 12 illustrates an example scene and demonstration. The state and action spaces, as well as initial states, are predefined in the scene. The program then generates HL subgoal states and LL actions required to accomplish the target task. The order of the 10 sub-tasks is interchangeable, provided they collectively achieve the target, which enhances the model’s generalization capability in real-world scenarios. Each sub-task has a maximum of 10 actions and will receive a “+1+1” reward when the sub-task is finished. Therefore, the planner should generate successful subgoals and actions while reducing the number of actions for each sub-task.
To diversify demonstrations, we randomized the initial and terminal states of objects, varied the sub-task sequences, and introduced redundant behaviors. This approach ensured suboptimal demonstrations with extended trajectories. Based on the average trajectory lengths in the demonstrations, tasks were categorized into Easy (50 steps), Medium (70 steps), and Hard (90 steps). Additionally, a subset of tasks with fixed terminal states was defined as “Single,” while the full dataset with randomized terminal states was labeled “Multi.” The dataset consisted of $10,000$ training samples and $1,000$ validation samples. Finally, we evaluated the planned tasks in the simulator by measuring the number of completed tasks and the steps required to execute them. We evaluate the planners’ performance in terms of the number of successfully finished sub-tasks and the normalized steps to finish each sub-task. The preferred optimal planning should complete more sub-tasks and the cumulated action steps should be minimized.
@@
Average
demo. length
Episode
length
Segment
interval
Planner horizon Diffusion steps
Single1 HL2 LL2 Single HL $L$
@@
We elaborate on implementing Bits Diffusion [18] in the task planning experiments. In Fig. 14, we first defined the whole state and action spaces, which were tokenized using the PreTrainedTokenizerFast in Huggingface (code). Then we converted the tokens as 8 Bits analog values. Since the state and action have two tokens, the diffusion model generates values with dimension “horizon $\times 16$”. In the sampling stage, the planners sampled values in continuous space. And we converted them with thresholding to binary values. Finally, the bits were digitalized and de-tokenized back to the texts for the following steps.